### v2

@@ 3 Quantum Computing Lab, Academic Computer Center Cyfronet AGH, Nawojki 11, Krakow, 30-950, Poland
Supplementary figures
Figure S.1: Hierarchical annealing in the Karate Club network for different resolution parameters. a Measure of the average overlap between the communities found by the Louvain and Hierarchical annealing algorithms (left axis, blue) and the number of communities (right axis, salmon and black) as a function of the resolution parameter $\gamma$. b Relative increase of the Hierarchical annealing measured w.r.t. the Louvain (top) and Leiden (bottom) solutions in a. Black, blue, and red symbols depict equal, better, and worse performance of the quantum method respectively. c Maximum modularity per resolution value (left axis, same legend as in a) and the total computing time per solution (right axis, green).
Figure S.2: Dendrogram of the Karate network [39]. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 50 and resolution = 1. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.3: Dendrogram of a clustered power-law network [35] of N=80 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 10 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.7563$. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.4: Dendrogram of a Barabasi-Albert network [34] of N=70 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 20 and resolution = 1. The Louvain and Leiden algorithms returned maximum modularities of $Q=0.0760$ and $Q=0.0698$ respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.5: Dendrogram of an Erdos-Renyi network [33] of N=70 nodes. Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 50 and resolution = 1. The Louvain and Leiden algorithms returned maximum modularities of $Q=0.1189$ and $Q=0.1021$ respectively. The left axis shows the modularity at each step of the process, while the right axis displays the corresponding increments. The bottom-most row represents the final output of the H. annealing algorithm.
Figure S.6: Dendrogram of a directed sclae-free network [36] of N=50 nodes Hierarchical structure found by the H. annealing algorithm. The hierarchy corresponds to the one unraveled after maximizing the modularity of the network as described in the main text; that is, num_runs = 50 and resolution = 1. Both the Louvain and Leiden algorithms returned a maximum modularity of $Q=0.31975$.
Figure S.7: Structural communities found at the early steps of the H. annealing process. We plot the community structure discovered during the 2nd division of the hierarchical method described in the main text. It is visible how the hierarchy carries concise anatomical information, thus being a valid way to inspect hidden structures within complex networks through quantum annealing optimization.
Figure S.8: Structural communities found by the Leiden and Louvain algorithms. The resemblance between these communities and the ones found by the annealing process is strikingly high (see main text). The modularity found by the Leiden algorithm was $Q=0.608$, and the modularity returned by the Louvain algorithm was $Q=0.611$.
Software specifics: Qommunity
Architecture
Figure S.9: Architecture of Qommunity library. The package makes use of the quantum computing library QHyper [65] to communicate with the quantum annealer using multiple instances of a quadratic binary optimization problem. The Qommunity library also allows the user to choose from several other algorithms and solvers described in the main text. Lastly, the modularity maximization happens at the IterativeSearcher class, where any solver and algorithm can be explicitly and automatically run an arbitrary number of times to produce a sufficiently big number of solutions to choose from.
@@
The Qommunity library was designed to integrate several methods for detecting communities in graphs in one place. The implemented methods rely on classical, quantum, and hybrid solutions. The idea was to simplify and standardize the interface for conducting experiments. Several components compartmentalize the architecture, offering user-friendly features: Samplers, Searchers, and IterativeSearchers (Fig. S.9).

### v3

@@ Acknowledgments
S3 Sampling the Hierarchical Annealing procedure a reasonable number of times
Because our algorithm has inherent stochasticity, the number of times we observe the maximum modularity value across repeated runs can be modeled as a discrete random variable. Let a success denote the event that a single run of the hierarchical annealing procedure achieves the maximum modularity value $\max Q$ (the global optimum for the problem under study). If we repeat the experiment independently $n$ times (i.e., $N_{runs}$), and if $p$ denotes the probability of observing this optimal value in a single run, then the number of successes $X$ follows a binomial distribution,
$X\sim\mathrm{Binomial}(n,p).$ (S3.1)
The probability of observing exactly $k$ successes is
$\begin{split}P(X=k)&=\binom{n}{k}p^{k}(1-p)^{n-k}\\ &=\frac{n!}{k!(n-k)!}p^{k}(1-p)^{n-k},\end{split}$ (S3.2)
for $k=0,1,\dots,n$. We are primarily interested in the probability of observing at least one success: achieving $\max Q$ at least once across the $n$ independent runs:
$\begin{split}P(X\geq 1)&=\sum_{r=1}^{n}P(X=r)\\ &=1-P(X=0)\\ &=1-(1-p)^{n}.\end{split}$ (S3.3)
The challenge is that the success probability $p$ is unknown in practice, and it may vary substantially depending on the structure and modularity landscape of the complex network under investigation. Consequently, before determining an appropriate number of repetitions $n$ for the hierarchical annealing procedure, we must form a reasonable estimate of $p$. Ideally, one could run the experiment an arbitrarily large number of times. However, each run is costly; for example, executing a single experiment on DWAVE consumes a non-negligible number of computing minutes. Although repeatedly running the experiment guarantees that we will eventually observe $\max Q$ even when $p$ is small (Fig. S2a; see also Eq. S3.3), doing so is highly inefficient for most plausible values of $p$. In practice, once we are only concerned with observing at least one success ($k\geq 1$), excessively repeating the experiment yields diminishing returns and substantial waste of computational resources.
S3.1 Estimating the probability of success
Suppose we repeat the hierarchical annealing procedure a total of num_runs times. Using Laplace’s principle, we can estimate the probability of observing a success by the empirical frequency with which the algorithm achieves $\max Q$:
$p\approx\frac{\text{\# times where }Q=\max Q}{N_{runs}}.$ (S3.4)
However, this estimate can be highly variable when the number of repetitions is small. In our case, with only $\texttt{num\_runs}=20$, the resulting estimate in Eq. S3.4 may be too noisy to rely on directly.
To obtain a more robust estimate of the probability of observing the maximum modularity value, we used a nonparametric bootstrap procedure. For each network, we treated the set of modularity values obtained from the original num_runs repetitions as the empirical distribution. We then generated bootstrap samples by resampling these values with replacement. Each bootstrap sample consisted of $N_{\mathrm{resamples}}$ draws, ranging from 10 to 1000 in increments of 10, to study the convergence of the boostrapped estimate of $p$. For all bootstrap samples, we computed the relative frequency with which the maximum modularity $\max Q$ was observed. Repeating this procedure $N_{\mathrm{experiments}}=500$ times for each value of $N_{\mathrm{resamples}}$ yielded a collection of estimates of the success probability. We recorded the mean of these bootstrap estimates to characterize the expected likelihood of observing a success $\mathcal{L}(\max Q)\to p$. This provided a non-parametric approximation of the underlying probability in Eq. S3.4.
Once we obtained a stable estimate of $p$ from the bootstrap procedure with $N_{\mathrm{resamples}}=1000$, we computed the theoretical probability of observing the maximum modularity at least once (Eq. S3.3). Across different networks, this probability remained above 50% for most cases and dropped to 40% in only two instances (Fig. S2b). These results confirm that even if the likelihood of observing $\max Q$ is relatively small (Fig. S2c), running the hierarchical annealing procedure 20 times provides a reasonable chance of achieving very high modularity values once.
Furthermore, to provide a transparent account of the behavior of the hierarchical annealing procedure, we report both $\min Q$ and $\max Q$ along with the corresponding bootstrapped likelihood estimates (Tables S9-S12). As expected, the frequency of observing $\max Q$ decreased as the number of nodes increased, but the bootstrap estimate of $p$ remained stable as $N_{\mathrm{resamples}}$ increased. Notably, the most likely value of $Q$ did not differ substantially from $\max Q$, reflecting both the high degeneracy of the modularity function and the very competitive performance of the recursive sampling performed by the DWAVE Advantage annealer.
S4 Computational complexity
The computational complexity of the recursive part of Algorithm 1 (see Main text) can be estimated using big-$\mathcal{O}$ notation, where $n$ indicates the size of the problem (i.e., number of nodes), $k$ indicates a given step, and $T(n_{k})$ depicts the time needed to perform the computation for $n_{k}$.
S4.1 Generalized modularity matrix
The modularity matrix B requires $\mathcal{O}(n^{2})$ computations, but it is calculated and stored only once at the start of the recursion. Then, it is read to obtain the generalized modularity matrix Bk in $\mathcal{O}(n)$ linear time, since a single summation over the nodes within a given community needs to be performed. Thus, in each level of the recursion, the computations involve linear operations and the realization of a quantum annealing (QA) step.
S4.2 Recursion tree
Several cases must be carefully considered. Moreover, the general case is better understood after analyzing the complexity of simpler scenarios. For this reason, we start with the worst-case scenario and build up to the most realistic approximation of the problem.
Additionally, as stated in the main text, the running time of the quantum annealing (QA) step can vary from polynomial [Farhi2001] to exponential times [Albash2018]. However, this is problem-specific and difficult to address analytically. For this reason, here we denote the complexity of a single quantum split as $f(n)$. Crucially, the final complexity and scalability of the algorithm will be modulated by the total QA running time, $\mathcal{K}_{n}$, but the functional dependence of the classical counterpart will not be affected by it.
Quadratic complexity
Although very unlikely, some networks could be recursively divided into two branches where one of the two can no longer be split in the next recursion steps. That is, $n_{k}=n_{k+1}+a\ \ \forall k$ and $a\in(0,n)$. In this case, the total time to divide the entire network into two communities can be written as
T(n)=T(na)+T(a)+n+f(n)+1,T(n)=T(n-a)+T(a)+n+f(n)+1,
and, subsequently,
T(na)=T(n2a)+na+f(na)+1,T(n-a)=T(n-2a)+n-a+f(n-a)+1,
$\ldots$
Crucially, the ”right” communities, always containing $a$ nodes, are not split again, thus no longer contributing to the tree. Then, the general running time for the $k$-th step is easily calculated by recursively substituting the different running times.
$\begin{split}T(n)&=T(n-ak)+\sum_{r=0}^{k-1}\left[1+n-ar+f(n-ar)\right]\\ &=T(n-ak)+k+nk-\frac{a}{2}(k-1)k+\sum_{r=0}^{k-1}f(n-ar)\end{split}$ (S4.1)
Assuming $n_{k}=n-ak\leq a$ as a stopping criterion yields $k=\frac{n}{a}-1>0$ number of steps. Therefore,
$\begin{split}T(n)&=T(a)+\frac{n}{a}-1+n\left(\frac{n}{a}-1\right)+\frac{a}{2}\left(\frac{n}{a}-2\right)\left(\frac{n}{a}-1\right)+\sum_{r=0}^{n/a-2}f(n-ar),\\ &\underset{n\gg 1}{\approx}\frac{3n^{2}}{2a}+\sum_{r=0}^{n/a}f(n-ar)\longrightarrow\mathcal{O}(n^{2}+\mathcal{K}_{n})\end{split}$ (S4.2)
where $T(a)$ is a constant, and $\mathcal{K}_{n}$ is the overall QA running time.
Unbalanced recursive trees
The quadratic instance of the recursion represents the worst-case scenario. In practice, however, each community found at step $k$ is typically subdivided in step $k+1$, producing two new communities of unequal size,
$n_{k}\frac{1}{a}\ \text{ and }\ n_{k}\left(1-\frac{1}{a}\right),$ (S4.3)
with $a\in(1,M)$ for some finite $M$. Notably, the stopping criterion is determined by the properties of the generalized modularity matrix, which differs for all discovered communities. Fortunately, due to the resolution limit, communities with fewer than $\sim\sqrt{m}$ nodes (modulated by $\gamma^{1/2}$ for arbitrary resolution parameters), where $m$ is the total number of edges, cannot be resolved [Fortunato2007, Kumpula2007]. This implies that the recursion will halt no later than at some minimum effective size $n_{0}\in[1,n-a]$. Then, considering the shrinking in Eq. S4.3 and setting $T(n_{0})=\mathcal{O}(1)$, the Akra-Bazzi theorem can be used to derive a tight bound on the algorithm’s complexity. Briefly, the theorem states that recurrences of the form
$T(n)=\sum_{i=1}^{s}a_{i}T(b_{i}n+h_{i}(n))+g(n),$ (S4.4)
have an asymptotic solution
$T(n)\in\Theta\left(n^{p}\left(1+\int_{n_{0}}^{n}\frac{g(u)}{u^{p+1}}du\right)\right),$ (S4.5)
where $p$ satisfies the characteristic equation
$\sum_{i=1}^{s}a_{i}b_{i}^{p}-1=0,$ (S4.6)
with $a_{i}>0$, $0<b_{i}<1$ for all $i=1,\ldots,s$, and $g(n)$ is a continuous, positive, and regular function for all sufficiently large $n$. Crucially, the QA complexity $f(n)$ seems exponential at worst [Albash2018], which still satisfies the conditions of regularity. The functions $h_{i}(n)$ represent small perturbations in the arguments of $T$, which, in our case, arise because they are integers rather than real numbers. In this case, $h_{i}(n)\in\mathcal{O}(1)$, since the rounding introduces a constant.
Applying the theorem to our case with $s=2$, $a_{1}=a_{2}=1$, $b_{1}=\frac{1}{a}$, and $b_{2}=1-\frac{1}{a}$ yields the initial recursion
$T(n)=T\left(\left[\frac{n}{a}\right]+c_{1}\right)+T\left(\left[n\left(1-\frac{1}{a}\right)\right]+c_{2}\right)+n+f(n),$
where $\left[x\right]$ is the integer function, and $c_{1},\ c_{2}$ are some unknown constants acting as the perturbations. Then, the corresponding characteristic equation
$\left(\frac{1}{a}\right)^{p}+\left(1-\frac{1}{a}\right)^{p}-1=0$
[…]
The asymptotic solution for Eq. S4.5 is then
$T(n)\in\mathcal{O}\left(n+n\log n+n\int_{n_{0}}^{n}\frac{f(u)}{u^{2}}du\right)\sim\mathcal{O}\left(n\log n+\mathcal{K}_{n}\right),$ (S4.7)
being
$\mathcal{K}_{n}\doteq n\int_{n_{0}}^{n}\frac{f(u)}{u^{2}}du$ (S4.8)
the total complexity derived from the repeated QA instances.
Unbalanced and dynamic recursive trees
There is another important consideration. The most realistic scenario corresponds to the following:
$n_{k}\frac{1}{a_{k}}\ \text{ and }\ n_{k}\left(1-\frac{1}{a_{k}}\right),$ (S4.9)
that is, the fraction of nodes assigned to each community changes along the hierarchy. However, this does not change the final result in Eqs. S4.7 and S4.8.
To understand this, we analyze the dynamics of the sequence $\{a_{k}\}_{k=0,...}$ representing the fraction of nodes assigned to each community at the $k$-th step. The elements $a_{k}$ can be modeled as uniform perturbations
$a_{k}=a+\delta_{k}+O(\delta_{k}^{2})$
where $a\in(1,M]$ and $\delta_{k}\sim U(-a,M)$ for some finite $M$ that depends on the average split size along the tree. The community sizes can then be expressed as,
$\begin{split}\frac{n_{k}}{a_{k}}&=\frac{n_{k}}{a}\left(\frac{1}{1+\delta_{k}/a}\right)=\frac{n_{k}}{a}\left(1-\frac{\delta_{k}}{a}+O(\delta_{k}^{2})\right)=\frac{n_{k}}{a}+h_{1}(n_{k})\\ n_{k}\left(1-\frac{1}{a_{k}}\right)&=n_{k}\left(1-\frac{1}{a}\frac{1}{(1+\delta_{k}/a)}\right)=n_{k}\left(1-\frac{1}{a}+\frac{\delta_{k}}{a^{2}}+O(\delta_{k}^{2})\right)=n_{k}\left(1-\frac{1}{a}\right)+h_{2}(n_{k})\end{split},$ (S4.10)
where $h_{1}(n_{k})$ and $h_{2}(n_{k})$ represent the perturbations in the sizes of the two communities, and both are of $\mathcal{O}(n_{k})$. These perturbations are bounded and decrease along the recursion. Therefore, even in the most realistic case of dynamic and unbalanced trees, the classical part of Algorithm 1 remains $\mathcal{O}(n\log n)$, and the quantum annealing complexity is still $\mathcal{K}_{n}$.
S4.3 Quantum annealing
We now evaluate the total complexity contributed by the quantum annealing steps, $\mathcal{K}_{n}$, under different assumptions for the function $f(n)$, which accounts for the cost of a single QA subproblem:
  1. I.
    If f(n)=cwith(0<c<M)f(n)=c\ $\text${with}\ (0<c<M), then $\mathcal{K}_{n}=O(1)$ and $T(n)\in\mathcal{O}(n\log n)$;
  2. II.
    if $f(n)\sim\log n$, then $\mathcal{K}_{n}=O(\log n)$ and $T(n)\in\mathcal{O}(n\log n+\log n)\sim\mathcal{O}(n\log n)$;
  3. III.
    if $f(n)\sim n$, then $\mathcal{K}_{n}=O(n\log n)$ and $T(n)\in\mathcal{O}(n\log n)$;
  4. IV.
    if $f(n)\sim n^{s},\ s>1$, then $\mathcal{K}_{n}=O(n^{s})$ and $T(n)\in\mathcal{O}(n\log n+n^{s})\sim\mathcal{O}(n^{s})$;
  5. V.
    and, if $f(n)\sim e^{n^{s}}\ \text{with}\ s\in\mathbb{R}^{+}$, then $\mathcal{K}_{n}$ has no closed-form expression, but the exponential complexity dominates, and $T(n)\sim\mathcal{O}(ne^{n^{s}})$.
One-hot encoding constraints
The community detection problem naturally lies in the discrete domain, where each node is assigned to one of the $|C|$ communities discovered. As discussed in the main text, QA requires the computational variables (i.e., the nodes) to be encoded in the binary domain (i.e., $x_{i}=\{0,1\}\ \forall i=1,\ldots,n$). A simple solution to this inconsistency is to use one-hot encoding, where each node’s community assignment is represented by a vector with a single 1 at the index of its assigned community and 0s elsewhere. That is,
$\mathbf{x}_{i}=(0,\ldots,1,\ldots,0),\ \forall i=1,\ldots,n.$
Since each node belongs to one and only one community, these one-hot vectors must satisfy the following property,
$\sum_{i=1}^{n}\left(\sum_{j=1}^{M}x_{ij}-1\right)^{2}=0,$
where $M$ is the dimension of the vectors. The constraint is imposed on the QUBO function via a Lagrange multiplier $\lambda$, whose optimal value is not known a priori. There is ongoing research on this topic [roch_2021, mayowa_2022]. However, in practice, $\lambda$ is often set through trial-and-error or grid search. Additionally, $M$ defines an upper bound on the number of detectable communities. Thus, even with an optimal $\lambda$, the choice of $M$ remains an open issue. We aim to understand how $M$ should scale with the number of nodes $n$. That is, if $M$ is appropriate for a network of size $n$, will it remain suitable for a network of size $2n$? And how will the computational complexity of the recursion be affected by this?
While in the extreme case $\gamma\to\infty$ modularity maximization would favor each node forming its own community (i.e., $M=n$), this limit is not realistic in practice [Kumpula2007]. The resolution limit inherent to the modularity function $Q$ implies that the number of communities that can be reliably detected is bounded above, with the exact maximum depending on network topology. This upper bound tends to scale between $\sqrt{n}$ and $n$ in sparsely and densely connected networks, respectively  [Fortunato2007, Kumpula2007]. With this in mind, revisiting the complexities of a QA step $f(n)$, it’s trivial to see that the only case where using a one-hot encoding scheme clearly favors the scalability is if the running time remains constant w.r.t. the size of the network (i.e., case I). Additionally, although using recursion theoretically degrades scalability when the annealing complexity is logarithmic, i.e., case II, the need to incorporate Lagrange multipliers in the one-hot encoding scheme may counteract this improvement. Even if we set $M=\sqrt{n}$,
  1. I.
    If f(n)=cwith(0<c<M)f(n)=c\ $\text${with}\ (0<c<M), the recursion is $\mathcal{O}(n\log n)$ while the one-hot encoding remains constant;
  2. II.
    if $f(n)\sim\log n$, the recursion is $\mathcal{O}(n\log n)$ while the one-hot encoding is $\mathcal{O}(\log n\sqrt{n})\sim\mathcal{O}(\log n)$
  3. III.
    if $f(n)\sim n$, the recursion is $\mathcal{O}(n\log n)$ while the one-encoding is $\mathcal{O}(n\sqrt{n})$;
  4. IV.
    if $f(n)\sim n^{s}\ \text{with}\ s>1$, the recursion is $\mathcal{O}(n^{s})$ while the one-hot encoding is $\mathcal{O}(n^{s}n^{s/2})$;
  5. V.
    and, if $f(n)\sim e^{n^{s}}\ \text{with}\ s\in\mathbb{R}^{+}$, the recursion is $\mathcal{O}(ne^{n^{s}})$ while the one-hot encoding is $\mathcal{O}(e^{n^{s}n^{s/2}})$.
Examples III to V might benefit from the explicit calculation when $n\to\infty$. Thus,
[…]
$\lim_{n\to\infty}\frac{n^{s}}{n^{s}n^{s/2}}=\lim_{n\to\infty}n^{-s/2}=0$
$\lim_{n\to\infty}\frac{ne^{n^{s}}}{e^{n^{s}n^{s/2}}}=\lim_{n\to\infty}ne^{-n^{s}(n^{s/2}-1)}=0$
favouring, in both cases, the recursion over an explicit one-hot encoding.
In the scenario where $M=n$, the complexity derived from $f(n^{2})$ scales even worse. In this aspect, using recursion and binary QA significantly improves the scalability of the algorithm, since it is a process that takes on average $n\log n$ time to complete. However, the practical speed-up stemming from the usage of quantum resources might largely depend on the running times of the QA processes themselves.
S5 QUBO and the dynamical range of the modularity
The $k$-th element $p_{k}$ of the process to optimize using QA is given by the elements of the generalized modularity matrix Bk (see Eqs. 10 and 11 in Methods),
$p_{k}=\frac{-1}{m}\sum_{ij\in k}B_{ij}^{\text{k}}x_{i}x_{j}.$ (S5.1)
The constant $m$, corresponding to the number of edges in the network, plays no role in the optimization, but it is kept here for completeness. To analyze the information (measured in bits) needed to encode the entries of the modularity matrix, we can compute the dynamic range [mucke_optimum-preserving_2024],
$\mathcal{DR}\doteq\ \log_{2}\left(\frac{\max D}{\min D}\right),$ (S5.2)
where $D$ is the set of absolute differences between all the elements of the (generalized) modularity matrix,
$D(\mathbf{B}^{\text{k}})\doteq\{|B_{ij}^{\text{k}}-B_{rs}^{\text{k}}|:\quad B_{ij}^{\text{k}},B_{rs}^{\text{k}}\in\mathbf{B}^{\text{k}},\quad|B_{ij}^{\text{k}}-B_{rs}^{\text{k}}|\geq\epsilon\}.$
The parameter $\epsilon$ controls for the numerical precision, and it was taken as $10^{-6}$. In principle, at each step of the recursion, the information decreases as the size of the communities diminishes and the ground state of the QUBO approaches the trivial solution. Because of this, we can treat the $k=0$ element as an upper bound, simplifying the numerical study, i.e., $D(\mathbf{B}^{\text{k}=0})=D(\mathbf{B})$.
To evaluate the sensitivity of Advantage to the $\mathcal{DR}$ and to create unrealistically harsh conditions, we simulated two unrealistic experiments. Initially, we generated chains of 3-cliques with 2, 6, and 20 3-cliques (i.e., 6,18, and 60 nodes, respectively). Then, two randomly selected edges were monotonously changed to (1,1), (10,.1), (50,.05), (100,.001), and the modularity was maximized with both Louvain and Hierarchical annealing (see Algorithm 1 in the Methods). In the second experiment, we generated 3 power-law clustered networks [Holme2002] of sizes $N=20,50,100$ with parameters $m=1$ and $p=.1$. Similarly, one randomly selected edge was monotonously changed to 1, 10, 50, 100, 500, 1000, 5000, and 10000. The performance of the Hierarchical annealing algorithm was compared to the Louvain in the same manner (Fig. S21).
S6 Embedding strategy, chain strength, and noise
S6.1 Setting the chain strength in the embedding
The modularity function $Q$ can be mapped onto a QUBO problem whose topology is determined by the entries of the modularity matrix $\mathbf{B}$. The solution is encoded in the maximization of the modularity,
$\hat{Q}\doteq\max_{\forall C\in\mathcal{C}}Q(C)=-\min_{\forall C\in\mathcal{C}}\bigl[-Q(C)\bigr].$ (S6.1)
In contrast, the quantum annealer (i.e., the QPU) evolves toward the minimum-energy configuration of the Hamiltonian of a distinct physical system, $H(\mathbf{J})$, which depends on the physical couplings $\mathbf{J}$ between qubits. The mapping from the QUBO formulation to the D-Wave annealer is achieved via a minor embedding procedure [Choi2008, choi_minor-embedding_2011]. This embedding defines the physical couplings between qubits $a$ and $b$ as a function of the logical couplings between nodes $i$ and $j$ in the network, such that $\mathbf{J}=f(\mathbf{B},k)$. Except in special cases where the logical and physical topologies are nearly identical, this mapping requires the introduction of chains of physical qubits to represent single logical variables. Within each chain, physical qubits are coupled by a chain strength $k$, which must be chosen to ensure a faithful representation of the logical problem while avoiding excessive distortion of the physical energy landscape [Grant2022]. In practice, the annealer seeks the ground state of $\mathcal{H}$ under the constraint that all qubits within a chain adopt the same state. When this constraint is violated, broken chains are resolved using a majority voting strategy, whereby the logical value is assigned according to the most common spin state along the chain.
To set the chain strength, we adopted D-Wave’s default strategy, which aims to mitigate inhomogeneities in the effective coupling strengths induced by the specific minor embedding [raymond2020improving]. This approach defines the chain strength as a function of both the logical problem and the topology of the physical connectivity graph:
$k^{*}=k_{0}\sqrt{\sigma^{2}N},$ (S6.2)
where $\sigma^{2}=\frac{2}{N(N-1)}\sum_{a<b}J_{ab}^{2}$ denotes the variance of the physical couplings, $N$ is the number of logical variables, and $k_{0}=\sqrt{2}$ was determined empirically [raymond2020improving]. Alternatively, the chain strength $k$ can be manually tuned by evaluating multiple values and selecting the one that yields the best empirical performance.
S6.2 Relationship between the chain strength and noise
Quantum annealers are intrinsically sensitive to noise, which in practice means that the device evolves toward the ground state of a perturbed Hamiltonian,
$H^{\delta}=H(\mathbf{J}+\delta\mathbf{J}),$ (S6.3)
where $\delta\mathbf{J}$ models deviations between the programmed and realized couplings. Under the integrated control error (ICE) model, these deviations are described as independent, zero-mean Gaussian perturbations with finite variance [yarkoni2022quantum]. Within this framework, a key observation is that the probability of chain breaks is related to the magnitude of the noise [Jeong2025], understood as the variance of $\delta\mathbf{J}$. Other factors at play include embedding-specific factors and energy scale separations (e.g., annealing temperature). Importantly, noise is itself a statistical abstraction: it provides an effective description of unknown and uncontrolled physical processes that give rise to discrepancies between the ideal results and empirical outcomes.
Consequently, an indirect strategy to assess the noise level of the QPU is to measure the fraction of broken chains, defined as the ratio between the number of chains that violate the embedding constraint and the total number of chains in a given embedded problem. This information can be directly extracted from the D-Wave samplers. For convenience, we define the chain break probability as the percentage of broken chains, $CBP=100\cdot CBF$, where $CBF$ denotes the chain break fraction. While this definition is operational, it is worth noting that $CBP$ and $CBF$ can be formally related to probabilistic measures through expectations under specific (and not always strictly satisfied) modeling assumptions [yamaji_development_2022].
S6.3 Experimental measurements of chain breaks and their relationship to annealing performance
To assess the performance of the annealing strategy proposed in this work and to evaluate whether noise constitutes a limiting factor, we systematically analyzed chain break behavior as a function of network size and the dispersion of outcomes across independent runs. Performance was quantified by the distance between individual solutions and the best result obtained, defined through the relative decrease in modularity as
$\text{Relative decrease}(N)\doteq\frac{Q(N)}{\max_{r\in N_{\mathrm{runs}}}Q(N)},$ (S6.4)
the relative distance was defined as
$100\cdot(1-\text{Relative decrease}).$
In addition, we measured the chain break fraction as a function of network size and computed statistical descriptors of the relative decrease in modularity across runs. Finally, to provide a quantitative assessment of the effectiveness of the default chain strength strategy, we manually tuned the chain strength $k$ for the largest network embeddable on the Pegasus topology ($N=169$) and compared the resulting performance against the default setting ($k=k^{*}$ in Eq. S6.2). The chain break fraction was defined as the maximum $CBF$ observed along the hierarchical tree, serving as a proxy for the most challenging subproblem associated with each instance. In practice, in almost all cases, this corresponded to the values measured at the highest level of the hierarchies (Fig. S22). For these experiments, we considered only Erdős–Renyi and power-law networks, as they represent two extremes of the same problem. The former corresponds to the canonical academic example of a random QUBO topology, whereas the latter represents a highly structured topology.
The relative decrease in modularity exhibited a clear degradation as a function of network size, accompanied by an increase in chain break fractions, suggesting an apparent global relationship between noise-related effects and solution quality (Fig. S23a). However, this trend did not persist when conditioning on individual network sizes (Fig. S23b). Specifically, for a fixed problem size, variations in chain break fractions across independent runs were not significantly associated with variations in solution quality (Pearson and Spearman $\rho$, $p$-values>0.05>0.05, exact tests). This indicates that the observed negative trends couldn’t be directly attributed to chain breaks at the level of individual instances. Rather, both increased chain break fractions and reduced solution stability emerged as concomitant consequences of scaling the logical problem size and its associated embedding complexity. In this sense, $CBF$s acted as a marker of problem difficulty under scaling of logical variables, rather than as a causal determinant of solution quality.
Indeed, the proportion of broken chains increased with the number of logical qubits (Fig. S23c), although it remained below 10% in the worst cases. Interpreting chain breaks as an indirect measure of noise, this result suggests that noise increased with problem size, as expected, but remained within acceptable levels throughout. Interestingly, this increase was significantly more pronounced for unstructured than for structured topologies (i.e., Erdős–Renyi and Powerlaw networks, respectively).
Concurrently, the chain strength increased with problem size (Fig. S23d), following a power-law scaling with an exponent of approximately $0.5$ for Erdős-Renyi networks and $\sim 0.06$ for Powerlaw networks. In the former case, this behavior closely matched the analytical prediction in Eq. S6.2, as the variance of physical couplings in random Gaussian systems is expected to remain constant [raymond2020improving]. In contrast, for Powerlaw networks, the variance of the couplings played a dominant role in determining the chain strength, with only a weak contribution arising from the increasing number of logical qubits. Lower chain strengths are expected to induce larger chain break fractions and, consequently, reduced resilience to noise. However, this behavior was not observed (Fig. S23c). This indicates that not only the number of logical qubits, but also the underlying topology of the problem, plays a critical role in determining appropriate chain strengths. As a result, the design of ad hoc or in-house strategies for setting the chain strength becomes challenging, and it is therefore preferable, in principle, to rely on the default and well-characterized strategy proposed by D-Wave [raymond2020improving].
These topological differences, as well as the different chain strength values, were reflected in the distribution of the solutions (Fig. S23e). In Erdős–Renyi networks, the standard deviation of the modularity values exhibited a discrete jump from zero to a relatively constant level, remaining stable as network size increased. In contrast, for Powerlaw networks, the variability of the solutions progressively increased with network size. These distinct patterns of variability were also mirrored in the average quality of the solutions. For Erdős–Renyi networks, the average distance to the best solution showed an abrupt increase and then stabilized at approximately 3% from the optimum. Conversely, in Powerlaw networks, the average solution quality deteriorated gradually to similar levels. Thus, the degradation in the average performance was driven by the increasing variability of the solutions, which broadly tracked the rise in chain break fractions, rather than purely by chain breaks.
Crucially, our primary interest was not the average modularity, but the highest. Accordingly, a more informative characterization of the solution landscape can be achieved by examining the upper percentiles of the modularity distribution. Since the best solution corresponds to the 100th percentile, the relevant question is how the high percentiles evolve as the variability of the solutions increases; i.e., how far do they move the best result. These percentiles remained largely constant for all problem sizes once the variability became non-zero (Fig. S23f). It is also worth noting that for the largest problem size, $N=169$, which corresponds to the practical limits of the current hardware, a ”second jump” was observed, analogous to the one occurring at $N=60$. This behavior suggests that the annealer retained its ability to explore high-quality solutions largely independent of the noise and concurrent chain breaks. The degradation in average modularity could instead be attributed to increased noise levels, which led to higher chain break fractions and, consequently, larger standard deviations in the distribution of solutions.
Lastly, we assessed the robustness of the default strategy for setting the chain strength by manually varying the chain strength $k$ for the largest network that could be embedded on Advantage ($N=169$). For simple networks, such as the Karate club graph, the solution was effectively independent of the chain strength (Fig. S23g). In contrast, for larger networks, the sensitivity of the solutions to the value of $k$ was evident (Fig. S23h). The default value $k^{*}$ consistently stayed within a range that ensured optimal performance, without compromising the annealer’s ability to reach the best possible solutions. This observation justifies the use of the default strategy. However, in scenarios where sufficient computational resources are available and maximal precision is required, exploring a neighborhood around the default value may result in marginal but relevant improvements.
In conclusion, although noise increased with problem size, the default chain strength strategy, accounting for both the topology and size of the QUBO, was effective in keeping its impact under control. Importantly, the presence of noise did not hinder the annealer’s ability to explore the solution landscape, allowing it to sample both lower- and higher-quality than average solutions. As a result, the annealer consistently achieved highly competitive performance when compared with state-of-the-art algorithms for modularity maximization. Using the default strategy to set the chain strength yields noise-resilient annealing strategies whose performance is ultimately governed by the problem size and the topology of the embedding.
S7 Software specifics: Qommunity
S7.1 Architecture
@@
The Qommunity library was designed to integrate several methods for detecting communities in graphs in one place. The implemented methods rely on classical, quantum, and hybrid solutions. The idea was to simplify and standardize the interface for conducting experiments. Several components compartmentalize the architecture, offering user-friendly features: Samplers, Searchers, and IterativeSearchers (Fig. S1).